a jet plane traveling at a constant goes 1200 mi with the wind , then turns around and travels for 1000 mi against the wind . if the speed of the wind is a constant 50 mph , and the total flight took 4 hours , find the speed of the plane.
Answer by mananth(16946)

plane speed x mph
wind speed 50 mph

Distance with wind 1200 miles
Distance against wind 1000 miles

speed with wind x + 50 mph
speed against x -50 mph
Total time = 4 hours
Time with current= 1200 /( x + 50 )
time against current 1000 / ( x -50 )

Time with current + time against = 4 hours

1200/( x+50)+1000/(x-50)=4
LCD =(x+50)*(x-50)
multiply the equation by the LCD
we get
1200*(x -50)+1000(x+50)=4
1200x-60000+1000x+50000 =4x^2
2200x-10000=4x^2-10000
4x^2-2200x+10000-10000 = 0
4x^2-2200x= 0
/ 4
X^2-550 x=0
x^2=550x
x=550 mph
m.ananth@hotmail.ca